OpenStudy (anonymous):

Solve for x. -3x - 8 + 6x = 7

OpenStudy (unklerhaukus):

combine like terms on the left

sam (.sam.):

-3x - 8 + 6x = 7 minus the 6x-3x 3x - 8 = 7 add 8 to both sides 3x - 8 +8 = 7 +8 3x=15 divide both sides by 3 x=5

OpenStudy (anonymous):

get x on one side! then divide